OpenStudy (poopsiedoodle):

How do you figure out what number 7 is 5/7s of?

OpenStudy (anonymous):

Divide 7 by 5/7

OpenStudy (anonymous):

7*(5/7) = 9(4/5) 9(4/5)*(5/7)=7

OpenStudy (poopsiedoodle):

luv u gurl

OpenStudy (anonymous):

haha np